# Free-Lunch Color-Texture Disentanglement for Stylized Image Generation

### v2

@@ Abstract
Recent advances in Text-to-Image (T2I) diffusion models have transformed image generation, enabling significant progress in stylized generation using only a few style reference images. However, current diffusion-based methods struggle with fine-grained style customization due to challenges in controlling multiple style attributes, such as color and texture. This paper introduces the first tuning-free approach to achieve free-lunch color-texture disentanglement in stylized T2I generation, addressing the need for independently controlled style elements for the Disentangled Stylized Image Generation (DisIG) problem. Our approach leverages the Image-Prompt Additivity property in the CLIP image embedding space to develop techniques for separating and extracting Color-Texture Embeddings (CTE) from individual color and texture reference images. To ensure that the color palette of the generated image aligns closely with the color reference, we apply a whitening and coloring transformation to enhance color consistency. Additionally, to prevent texture loss due to the signal-leak bias inherent in diffusion training, we introduce a noise term that preserves textural fidelity during the Regularized Whitening and Coloring Transformation (RegWCT). Through these methods, our Style Attributes Disentanglement approach (SADis) delivers a more precise and customizable solution for stylized image generation. Experiments on images from the WikiArt and StyleDrop datasets demonstrate that, both qualitatively and quantitatively, SADis surpasses state-of-the-art stylization methods in the DisIG task. Code will be released at here.
{strip}[Uncaptioned image]
Figure 1: Stylized images generated by our training-free method, SADis. (Up) As shown in the first two rows, it enables disentangled control over color and style attributes in text-to-image diffusion models using separate image prompts. This approach offers creators enhanced color control, including the use of color palettes as in the last two columns. (Down) SADis also enables real-image stylization by incorporating a content image as an additional condition via ControlNet. Furthermore, it extends to color-only stylized generation and material transfer for more flexible image generation.
@@ 1 Introduction
Stylized Image Generation [12, 30, 36], also termed style transfer in classical research, aims to transfer a style from a reference image to a target image. This field has evolved through several technological paradigms, beginning with CNN-based feature manipulation [20, 65, 36], advancing to attention mechanisms [4, 74, 46], and further developing through GAN-based image translation [37, 81, 72]. A significant advancement came with multimodal CLIP guidance [18, 40, 52], which bridged the gap between visual and textual representations. This breakthrough enabled a novel approach where style transfer could be guided by textual descriptions rather than being limited to reference images [18, 41, 35]. The field underwent another transformation with the emergence of Text-to-Image (T2I) diffusion models [56, 59, 57], which demonstrated remarkable capabilities in personalized image generation [58, 19, 49, 7].
@@
T2I diffusion models catalyzed new developments in stylized image generation [54, 67, 62] — a specialized paradigm that focuses on creating new images that incorporate specific visual characteristics from reference styles, rather than simply transferring style between existing images. Initially, many approaches [19, 2] relied on extensive fine-tuning using datasets of similarly styled images. However, this requirement proved impractical in real-world scenarios, where collecting cohesive style-specific datasets is often challenging. Addressing these limitations, recent research has focused on developing tuning-free (free-lunch) methods [75, 54, 38]. These approaches eliminate the need for costly retraining while maintaining efficient style integration capabilities. Despite their methodological differences, both tuning-based and tuning-free approaches share a common characteristic: they transfer entangled style representations during the image generation process, simultaneously incorporating both color and texture elements from the style images into the final output.
@@
Although existing stylized image generation methods offer promising advancements, for content creators, these methods have critical limitations. A key challenge is the lack of granular control over style elements. Content creators often need to adopt specific aspects of a reference style while preserving elements of their original vision [14, 27]. Color palettes, in particular, play a crucial role in this process — they are often meticulously crafted to evoke specific emotions or maintain brand consistency. Creators may wish to preserve these while adopting textural [^1] In this paper, we use the term texture to refer to the arrangement, repetition, and local patterns of visual elements such as shapes, edges, intensities, and gradients and excluding color aspects. elements from other reference styles. However, current approaches force an all-or-nothing choice, where accepting a reference style means incorporating all its visual (texture and color) characteristics simultaneously, seriously limiting artistic freedom and practical applicability.
@@
To address these limitations, we introduce the problem of Disentangled Stylized Image Generation (DisIG), which aims to decompose reference styles into independently controllable attributes like color, texture, and semantic content. This formulation enables artists to selectively transfer specific style elements while maintaining control over the content using text prompts. One potential workaround for DisIG using existing methods is to specify desired style attributes through text prompts. However, this approach requires extensive prompt engineering [48, 76, 71] expertise and considerable trial-and-error to achieve results that can be obtained more intuitively through image prompts. Even with significant effort in crafting precise textual descriptions, text prompts often fall short of capturing the nuanced characteristics of the target style. This limitation stems from the inherent expressiveness gap between text and visual information — image prompts inherently contain richer semantic information, including style attributes such as color and texture, than text prompt descriptions, that nuanced details lead to enhance generation quality as shown in Fig. 2.
@@
In this work, we propose what we believe to be the first DisIG method that enables independent control of color and texture through separate reference images, without any training. To achieve this, we first analyze the additivity property — a characteristic previously studied in textual spaces [51, 6, 33] — and demonstrate that it also holds true in the image prompt space. Then, we work with each style component separately: we isolate the color representation through feature subtraction between the original color image and its grayscale equivalent, while deriving the texture representation using a grayscale version of the texture reference image and SVD rescaling. Finally, to ensure accurate color matching while preserving textural details, we introduce RegWCT, an enhanced Whitening and Coloring Transformation with noise regularization that counteracts the signal-leak bias [16, 80] common in diffusion models.
@@
In experiments, we use images from the WikiArt [64] and StyleDrop [62] datasets and evaluate our approach based on the SDXL model [53]. By comparing against several state-of-the-art stylization methods, SADis consistently outperforms these baselines over both qualitative and quantitative results, particularly in achieving accurate color and texture expression for the stylized T2I generation. In summary, our contributions are as follows:
@@
  • We are the first to identify and leverage the image-prompt additivity property in the CLIP image encoder, showing how image representations can be decomposed and combined similarly to text embeddings.
@@
  • Through comprehensive qualitative and quantitative evaluations in standard style transfer benchmarks, SADis consistently outperforms state-of-the-art baselines, especially in realistic T2I generation under the DisIG scenario.
Figure 2: Compared to existing stylization methods (InstantStyle and DEADiff), our SADis offers flexibility in style control by enabling the use of separate color and texture reference images [^3] Here we use the prompt “A ballerina” as an example. The prompt used in InstantStyle and DEADiff is a detailed description of the color reference image generated by GPT-4o [1]..
@@ 2 Related Work
Stylized image generation [30, 12, 37, 81], also related to classical style transfer, aims to generate images with the artistic style of a reference image. Early research focused on statistical feature manipulation to transfer artistic styles from reference images. Gatys et al. [20] pioneered this direction by using covariance matrices for style representation. Following works like AdaIN [36] improved efficiency by transferring feature statistics between style and content feature maps, while WCT [44] introduced whitening and coloring transformations to match covariance matrices. Recent architectural innovations have enhanced stylization capabilities. A significant paradigm shift occurred with the introduction of CLIP [55], enabling connections between text and image representations in a shared embedding space. This advancement spawned methods like ClipStyler [41], which combines global and patch-level CLIP losses for high-resolution stylization.
Figure 3: Our method, SADis, begins with color-texture extraction (CTE), leveraging the Image-Prompt Additivity property, which is verified in this paper for the first time. After obtaining the disentangled color and texture embeddings from the reference images, we incorporate these embeddings into the style cross-attention layer of the SDXL model. Subsequently, we refine the latent variable $z_{t-1}$ at each inference step with our proposed RegWCT transformation, aligning color palettes precisely while retaining essential texture details.
@@
Nowadays, T2I Diffusion models [61, 29, 9] have emerged as the new state-of-the-art for text-to-image generation [8, 45, 33, 47]. Then the existing T2I stylization methods [23, 60, 13, 78, 66, 22, 79] achieve stylized image generation via fine-tuning generative models on few style reference images. However, this process is time-consuming and struggles to generalize to real-world scenarios where gathering a suitable subset of shared-style images can be difficult. To address these limitations, interest in tuning-free (free-lunch) approaches for stylized image generation has grown [11, 15, 70, 69, 26, 24]. These methods introduce lightweight adapters that extract style information from reference images and inject it into the diffusion process via self-attention or cross-attention layers. Representative examples include IP-Adapter [75] and Style-Adapter [69], which employ a decoupled cross-attention mechanism that separates cross-attention layers for handling text and image features independently. DEADiff [54] introduces a different approach by focusing on extracting disentangled representations of content and style using paired datasets.
@@
Although existing stylized image generation methods show promise, they lack granular control over individual style elements - a crucial need for content creators who often want to selectively adopt specific aspects of reference styles. This introduces a challenge we term the Disentangled Stylized Image Generation (DisIG). Focusing on color and texture as the most significant style attributes, this paper introduces the first free-lunch approach to color-texture disentanglement for stylized image generation.
@@ 3.1 Preliminaries
Problem Setup. Unlike classical stylized image generation [44, 10, 40], which derives style information from a single reference image, Disentangled Stylized Image Generation (DisIG) leverages two distinct reference images: a color reference image $\mathcal{I}_{clr}$, and a texture reference image $\mathcal{I}_{tx}$. The color reference image could either be a normal RGB image or a color palette. Given an input prompt $\mathcal{P}$ for the T2I model, our goal is to generate an image aligning with the prompt while exhibiting a color palette similar to $\mathcal{I}_{clr}$ and texture from $\mathcal{I}_{tx}$.
@@
T2I Diffusion Models. We built on the SDXL [53] model, consisting of two primary components: an autoencoder and a diffusion model $\epsilon_{\theta}(z_{t},t,\tau_{\xi}(\mathcal{P}))$, where $\epsilon_{\theta}$ is a UNet, conditioning a latent input $z_{t}$, a timestep $t\sim\text{U}(1,T)$, and a text embedding $\tau_{\xi}(\mathcal{P})$. More specifically, text-guided diffusion models generate an image from the textual condition as $\mathcal{C}_{text}=\tau_{\xi}(\mathcal{P})$, where $\tau_{\xi}$ is the CLIP text encoder [55] [^4] SDXL uses two text encoders and concatenate the embeddings.. The cross-attention map is derived from $\epsilon_{\theta}(z_{t},t,\mathcal{C}_{text})$. After predicting the noise, diffusion schedulers [63, 50] are used to predict the latent $z_{t-1}$. Here we simplify the $z_{t-1}$ inference process as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text})$.
@@
IP-Adapter. Building on T2I diffusion models, the IP-Adapter [75] introduces additional controllability by conditioning the T2I model on a conditional image $\mathcal{I}_{ip}$. Practically, this involves leveraging a pre-trained T2I diffusion model and incorporating a cross-attention layer to the (projected) image condition following each text-prompt conditioning layer. The conditional image is encoded in the low-dimensional CLIP image embedding space [55] to capture high-level semantic information. By denoting the CLIP image encoder as $\tau_{\phi}$ and IP-Adapter projection as $\mathbf{IP}$, this process is adding a new image condition $\mathcal{C}_{img}=\mathbf{IP}(\tau_{\phi}(\mathcal{I}_{ip}))$ to the T2I model as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathcal{C}_{img})$.
Figure 4: By subtracting the grayscale image embeddings from the color image embeddings, we can effectively remove the texture information and isolate the color information. Combined with the prompt “A photo of a bird” we can generate images in the same color schemes as the input images.

### v3

@@ Abstract
Recent advances in Text-to-Image (T2I) diffusion models have transformed image generation, enabling significant progress in stylized generation using only a few style reference images. However, current diffusion-based methods struggle with fine-grained style customization due to challenges in controlling multiple style attributes, such as color and texture. This paper introduces the first tuning-free approach to achieve free-lunch color-texture disentanglement in stylized T2I generation, addressing the need for independently controlled style elements for the Disentangled Stylized Image Generation (DisIG) problem. Our approach leverages the Image-Prompt Additivity property in the CLIP image embedding space to develop techniques for separating and extracting Color-Texture Embeddings (CTE) from individual color and texture reference images. To ensure that the color palette of the generated image aligns closely with the color reference, we apply a whitening and coloring transformation to enhance color consistency. Additionally, to prevent texture loss due to the signal-leak bias inherent in diffusion training, we introduce a noise term that preserves textural fidelity during the Regularized Whitening and Coloring Transformation (RegWCT). Through these methods, our Style Attributes Disentanglement approach (SADis) delivers a more precise and customizable solution for stylized image generation. Experiments on images from the WikiArt and StyleDrop datasets demonstrate that, both qualitatively and quantitatively, SADis surpasses state-of-the-art stylization methods in the DisIG task.
Figure 1: Stylized images generated by our training-free method, SADis. (Up) Example of texture (left) and color (above) conditioned text-to-image (T2I) generation. This approach offers creators enhanced color control, including the use of color palettes as shown in the last two columns. (Down) Example with conditioning on content image (left) and color (above), showing it extends to color-only stylized image generation.
@@ 1 Introduction
Stylized Image Generation deng2022stytr2 ; hong2023aespa_net ; huang2017AdaIN aims to transfer a style from a reference image to a target image. This field has evolved through several technological paradigms, beginning with CNN-based feature manipulation gatys2016image_style_transfer ; tumanyan2022splicing ; huang2017AdaIN , advancing to attention mechanisms alaluf2024cross_image_attn ; yao2019attention_stroke ; liu2021adaattn , and further developing through GAN-based image translation isola2017image_pix2pix ; zhu2017unpaired_cyclegan ; wu2021stylealign . A significant advancement came with multimodal CLIP guidance gal2022stylegan_nada ; kim2021diffusionclip ; patashnik2021styleclip , which bridged the gap between visual and textual representations. This breakthrough enabled a novel approach where style transfer could be guided by textual descriptions rather than being limited to reference images gal2022stylegan_nada ; kwon2021clipstyler ; huang2024diffstyler . The field underwent another transformation with the emergence of Text-to-Image (T2I) diffusion models ramesh2022dalle2 ; saharia2022photorealistic ; Rombach_2022_CVPR_stablediffusion , which demonstrated remarkable capabilities in personalized image generation ruiz2022dreambooth ; textual_inversion ; Cones2023 ; butt2025colorpeel .
@@
T2I diffusion models catalyzed new developments in stylized image generation qi2024deadiff ; wang2024instantstyle ; sohn2023styledrop — a specialized paradigm that focuses on creating new images that incorporate specific visual characteristics from reference styles, rather than simply transferring style between existing images. Initially, many approaches textual_inversion ; agarwal2023image_matte relied on extensive fine-tuning using datasets of similarly styled images. However, this requirement proved impractical in real-world scenarios, where collecting cohesive style-specific datasets is often challenging. Addressing these limitations, recent research has focused on developing tuning-free (free-lunch) methods ye2023ip-adapter ; qi2024deadiff ; jeong2024visual_style_prompting . These approaches eliminate the need for costly retraining while maintaining efficient style integration capabilities. Despite their methodological differences, both tuning-based and tuning-free approaches share a common characteristic: they transfer entangled style representations during the image generation process, simultaneously incorporating both color and texture elements from the style images into the final output.
@@
Although existing stylized image generation methods offer promising advancements, for content creators, these methods have critical limitations. A key challenge is the lack of granular control over style elements. Content creators often need to adopt specific aspects of a reference style while preserving elements of their original vision epstein2023art ; hertzmann2022toward . Color palettes, in particular, play a crucial role in this process — they are often meticulously crafted to evoke specific emotions or maintain brand consistency. Creators may wish to preserve these while adopting textural [^1] In this paper, we use the term texture to refer to the arrangement, repetition, and local patterns of visual elements such as shapes, edges, intensities, and gradients and excluding color aspects. elements from other reference styles. However, current approaches force an all-or-nothing choice, where accepting a reference style means incorporating all its visual (texture and color) characteristics simultaneously, seriously limiting artistic freedom and practical applicability.
@@
To address these limitations, we introduce the problem of Disentangled Stylized Image Generation (DisIG), which aims to decompose reference styles into independently controllable attributes like color, texture, and semantic content. This formulation enables artists to selectively transfer specific style elements while maintaining control over the content using text prompts. One potential workaround for DisIG using existing methods is to specify desired style attributes through text prompts. However, this approach requires extensive prompt engineering liu2022design_guidelines ; yu2024seek_for_incantations ; witteveen2022investigating_prompt_engineer expertise and considerable trial-and-error to achieve results that can be obtained more intuitively through image prompts. Even with significant effort in crafting precise textual descriptions, text prompts often fall short of capturing the nuanced characteristics of the target style. This limitation stems from the inherent expressiveness gap between text and visual information — compared with text prompts, image prompts inherently contain more fine-grained semantic information that is difficult to describe accurately in text, including style attributes such as color and texture. Such nuanced details contribute to improved generation quality (as shown in Fig.˜2-2nd row.).
Figure 2: Compared to SOTA stylization methods, SADis enables more precise and fine-grained control of color and texture by leveraging style element images, which allows for more accurate specification of visual attributes [^3] Here we use the prompt “A ballerina” as an example. The prompt used in InstantStyle and DEADiff is a detailed description of the color reference image generated by GPT-4o achiam2023chatgpt4 ..
@@
In this work, we propose what we believe to be the first DisIG method that enables independent control of color and texture through separate reference images, without any training. Such application and performance can be seen from Fig.˜1. To achieve this, we first analyze the additivity property — a characteristic previously studied in textual spaces NIPS2013_wordembedding ; manuel2023sega ; hu2024tome — and demonstrate that it also holds true in the image prompt space. Then, we work with each style component separately: we isolate the color representation through feature subtraction between the original color image and its grayscale equivalent, while deriving the texture representation using a grayscale version of the texture reference image and SVD rescaling. Finally, to ensure accurate color matching while preserving textures, we introduce RegWCT, an enhanced whitening and coloring transform with noise regularization, which counteracts the signal-leak bias everaert2024_signal_leak_bias ; zhang2024real_bias in diffusion models.
@@
In experiments, we use images from the WikiArt artgan2018_wikiart and StyleDrop sohn2023styledrop datasets and evaluate our approach based on the SDXL model podell2023sdxl . By comparing against several state-of-the-art stylization methods, SADis consistently outperforms these baselines over both qualitative and quantitative results, particularly in achieving accurate color and texture expression for the stylized T2I generation. In summary, our contributions are as follows:
@@
  • We are the first to identify and use the image-prompt additivity property in the CLIP image encoder, showing how image features can be decomposed and combined similarly to text embeddings.
@@
  • Through thorough qualitative and quantitative evaluations in standard style transfer benchmarks, SADis consistently outperforms existing baselines in realistic generation under the DisIG scenario.
@@ 2 Related Work
Stylized image generation hong2023aespa_net ; deng2022stytr2 ; isola2017image_pix2pix ; zhu2017unpaired_cyclegan , also related to classical style transfer, aims to generate images with the artistic style of a reference image. Early research focused on statistical feature manipulation to transfer artistic styles from reference images. Gatys et al. gatys2016image_style_transfer pioneered this direction by using covariance matrices for style representation. Following works like AdaIN huang2017AdaIN improved efficiency by transferring feature statistics between style and content feature maps, while WCT li2017universal introduced whitening and coloring transformations to match covariance matrices. Recent architectural innovations have enhanced stylization capabilities.
A significant paradigm shift occurred with the introduction of CLIP radford2021clip , enabling connections between text and image representations in a shared embedding space. This advancement spawned methods like ClipStyler kwon2021clipstyler to combine global and patch-level CLIP losses for image stylization.
@@
Nowadays, T2I Diffusion models (deepfloyd, ; ho2022imagen, ; chen2023pixartalpha, ) have emerged as the new state-of-the-art models for text-to-image generation. Then the existing T2I stylization methods han2023svdiff ; shi2023instantbooth ; dong2022dreamartist ; zhang2023prospect ; voynov2023ETI ; hiper2023 ; zhang2023InST ; kai2023DPL ; li2023stylediffusion achieve stylized image generation via fine-tuning generative models on few style reference images. However, this process is time-consuming and struggles to generalize to real-world scenarios where gathering a suitable subset of shared-style images can be difficult. To address these limitations, interest in tuning-free (free-lunch) approaches for stylized image generation has grown chung2024styleID ; Everaert2023_diff_instyle ; wang2023stylediffusion ; wang2023styleadapter ; Hertz2024_style_aligned ; he2024freestyle . These methods introduce lightweight adapters that extract style information from reference images and inject it into the diffusion process via self-attention or cross-attention layers. Representative examples include IP-Adapter ye2023ip-adapter and Style-Adapter wang2023styleadapter , which employ a decoupled cross-attention mechanism that separates cross-attention layers for handling text and image features independently. DEADiff qi2024deadiff introduces a different approach by focusing on extracting disentangled representations of content and style using paired datasets.
Figure 3: Our method, SADis, begins with color-texture extraction (CTE), leveraging the Image-Prompt Additivity property, which is verified in this paper for the first time. The color embedding $\mathbf{Emb}_{clr}$ is obtained by exploiting the Image-Prompt Additivity property, while the texture embedding $\mathbf{Emb}_{tx}$ is extracted via a singular value decomposition (SVD) operation. Afterwards, we incorporate these embeddings into the style cross-attention layer of the SDXL model. Subsequently, we refine the latent variable $z_{t-1}$ at each inference step with our proposed RegWCT transformation, aligning color palettes precisely while retaining essential texture details.
@@
Although existing stylization methods show promise, they lack granular control over individual style elements - a crucial need for content creators who often want to selectively adopt specific aspects of reference styles. This introduces a challenge we term the Disentangled Stylized Image Generation (DisIG). Focusing on color and texture as the most significant style attributes, this paper introduces the first free-lunch approach to color-texture disentanglement for stylized image generation.
@@ 3.1 Preliminaries
T2I Diffusion Models. We built on the SDXL (podell2023sdxl, ) model, consisting of two primary components: an autoencoder and a diffusion model $\epsilon_{\theta}(z_{t},t,\tau_{\xi}(\mathcal{P}))$, where $\epsilon_{\theta}$ is a UNet, conditioning a latent input $z_{t}$, a timestep $t\sim\text{U}(1,T)$, and a text embedding $\tau_{\xi}(\mathcal{P})$. More specifically, text-guided diffusion models generate an image from the textual condition as $\mathcal{C}_{text}=\tau_{\xi}(\mathcal{P})$, where $\tau_{\xi}$ is the CLIP text encoder (radford2021clip, ) [^4] SDXL uses two text encoders and concatenate the embeddings.. The cross-attention map is derived from $\epsilon_{\theta}(z_{t},t,\mathcal{C}_{text})$. After predicting the noise, diffusion schedulers song2021ddim ; lu2022dpm are used to predict the latent which we simplify as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text})$.
@@
IP-Adapter. Building on T2I diffusion models, the IP-Adapter ye2023ip-adapter introduces additional controllability by conditioning the T2I model on a conditional image $\mathcal{I}_{ip}$. Practically, this involves leveraging a pre-trained T2I diffusion model and incorporating a cross-attention layer to the (projected) image condition following each text-prompt conditioning layer. The conditional image is encoded in the low-dimensional CLIP image embedding space radford2021clip to capture high-level semantic information. By denoting the CLIP image encoder as $\tau_{\phi}$ and IP-Adapter projection as $\mathbf{IP}$, this process is adding a new image condition $\mathcal{C}_{img}=\mathbf{IP}(\tau_{\phi}(\mathcal{I}_{ip}))$ to the T2I model as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathcal{C}_{img})$.
3.2 Disentangled Stylized Image Generation
In a typical artistic workflow, creators draw inspiration from multiple sources to produce their final results. They might select a color palette from a sunset photograph, adopt textural patterns from a baroque painting, and incorporate structural elements from architectural blueprints—all while maintaining precise control over which elements they adopt from each reference. We term this ideal multi-target style transfer problem as Disentangled Stylized Image Generation (DisIG).
This multi-source approach to artistic creation (DisIG) contrast to current computational stylization methods li2017universal ; cho2019i2i_wct ; kim2021diffusionclip , which typically extract and apply a bundled representation of “style” from a single reference image. To address the limitation, we first analyze how color and texture are entangled in existing stylization approaches. We focus specifically on these two attributes because they represent fundamental, measurable components of visual style that artists frequently manipulate independently.
Figure 4: a) Image-prompt additivity: Image encoder latent space enables color extraction via grayscale subtraction. This color information can then be added to latent representations of other grey-scale images; b) By subtracting the grayscale image embeddings from the color image embeddings, we can effectively remove the texture information and isolate the color information.
Problem Setup. We formally define the DisIG task as follows: Given a text prompt $\mathcal{P}$, a color reference image $\mathcal{I}_{clr}$, and a texture reference image $\mathcal{I}_{tx}$, our goal is to generate an image that semantically aligns with $\mathcal{P}$ while independently adopting the color distribution of $\mathcal{I}_{clr}$ and the textural characteristics of $\mathcal{I}_{tx}$. The color reference $\mathcal{I}_{clr}$ can be either a standard RGB image or a discrete color palette. Unlike traditional style transfer that applies a holistic "style" from a single reference, our formulation enables precise, attribute-specific control by explicitly disentangling and separately transferring color and texture components.